OpenStudy (anonymous):

Find the perimeter of a rectangle with l = 1.2 m and w = 0.6 m A. 0.72 m B.1.8 m C.3.6 m D. 7.2 m

OpenStudy (anonymous):

c

OpenStudy (a1234):

1.2 + 1.2 = 2.4 0.6 + 0.6 = 1.2 2.4 + 1.2 = 3.6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem